# Access Hoare Logic

### v2

@@ Abstract.
Following Hoare’s seminal invention, later called Hoare logic, to reason about correctness of computer programs, we advocate a related but fundamentally different approach to reason about access security of computer programs such as access control. We define the formalism, which we denote access Hoare logic, and present examples which demonstrate its usefulness and fundamental difference to Hoare logic. We prove soundness and completeness of access Hoare logic, and provide a link between access Hoare logic and standard Hoare logic.
@@
$c$s: Theory of computation Hoare logicccs: Theory of computation Logic and verificationccs: Software and its engineering Software verificationccs: Security and privacy Logic and verificationccs: Security and privacy Software security engineering
@@ 1. Introduction
The correctness of programs has been a concern early on in the development of computers. The seminal contribution by Tony Hoare (hoare:69) allows to reason about programs instruction by instruction using pre- and postconditions. It allows to ascertain that if the precondition is satisfied before execution, and program execution terminates, then the postcondition will be satisfied.
@@
This form of ascertaining the correctness of programs has been and still is very successful. It is at the heart of several tools that are used in academia and industry to specify and verify software systems, such as SPARK (SPARKAda:Homepage) and Dafny (dafny:Homepage).
@@
However, there are security properties that do not naturally fall within the framework of Hoare logic. Access control is the property of restricting access to a resource. It is a well-established area of computer security that includes identification, authorization, authentication, access approval, and auditing (gollmann:2011), where access approval is the process of deciding whether to grant or reject an access request. Access control has a broad range of applications, ranging from large distributed systems to local systems such as electronic keys. More recently, it gained importance in the context of blockchain and cryptocurrencies: Smart contracts are often written to control access to digital assets, such as cryptocurrencies, tokens, or other forms of digital assets managed by blockchain systems, thereby combining several properties of general access control. We will use the term access security to speak about such properties in general.
@@
For a program to be access secure, we need to guarantee that the final state (e.g., access to a hotel room via electronic keys) will only be granted if a given set of conditions has been met before execution (e.g., possession of valid access codes). That is, for access security we are interested in the reverse direction of Hoare logic, which we will denote access Hoare logic (aHl), where we will reason from postconditions to preconditions. While normal Hoare logic expresses a precondition to be sufficient for the postcondition to hold after executing a program, access Hoare logic requires the precondition to be necessary. More precisely, access Hoare logic expresses that if the program starting in state $s$ terminates in a state $s^{\prime}$ which fulfills the postcondition, then it is necessary that state $s$ fulfills the precondition.
Another form of reversing Hoare logic has been proposed in the literature (vriesKoutavas:ReverseHoareLogic) and further developed to deal with incorrectness of programs (OHearn:IncorrectnessLogic). That formalism expresses that for each final state $s^{\prime}$ satisfying the postcondition there is a state $s$ satisfying the precondition such that the program, when started in state $s$, can terminate in state $s^{\prime}$. This is a different perspective, which is incompatible with our approach, to the best of our knowledge.
@@
In this brief note, we define access Hoare triples and access Hoare logic and use it to describe access security for examples. We show that access Hoare logic is sound and complete for a standard, general interpretation of access Hoare triples. We provide a link between Hoare logic and access Hoare logic in that weakest preconditions for usual Hoare logic correspond to strongest preconditions for access Hoare logic on states where the program terminates.
@@
Alhabadi et al. (alhabadi:etal:2022) have recently discovered that weakest preconditions for Hoare logic can be used to express access security for Bitcoin scripts. Our observation here expands on those insights and provides a reason why that is the case.
@@ 2. Examples
We demonstrate our point of view by considering three examples: The first involves electronic keys, the second owning and transferring Bitcoins, and the third granting access depending on the presence of a key in a list of keys.
@@ 2.1. Access Security for Electronic Keys
Our first example is a digital key system in a hotel that uses cards to open doors, which was specified and verified by Jackson (jackson2012:softwareAbstraction:2ndEd) using Alloy, and by Nipkow (Nipkow-VerficationHotelKeySystem) using Isabelle/HOL. In this system, the battery-powered door locks are not connected to a network. Instead, their state is controlled by cards that are inserted into them, or, in wireless systems, moved close to them. A lock holds an internal electronic key $\mathtt{dk}$ which controls its state. A card holds two electronic keys, $\mathtt{ck1}$ for the first key and $\mathtt{ck2}$ for the second. When a new guest receives a card (for example, from the hotel reception), it will contain the key corresponding to a previous guest as $\mathtt{ck1}$, and a new fresh key as $\mathtt{ck2}$. An access control program for the door lock suitable for use in such an electronic hotel key system should give access to the room in an access-secure way, by setting a variable $\mathtt{acc}$ to true: When the guest uses their card for the first time, access will be granted if the previous guest’s key is stored as the door key ($\mathtt{dk}{}=\mathtt{ck1}{}$), in which case the door key is updated to the new key $\mathtt{ck2}$, preventing the previous guest from gaining access to the room thereafter. The next time the new guest swipes their card, access will be granted because the second key on their card, $\mathtt{ck2}$, now matches the door key ($\mathtt{dk}{}=\mathtt{ck2}{}$).
For this example, a preconditon for verifying access security is given by $\mathtt{ck1}=\mathtt{dk}\lor\mathtt{ck2}=\mathtt{dk}$, expressing that either of the first key or the second key is the door key. A postcondition is $\mathtt{acc}=\mathtt{true}$, expressing that access is granted.
@@
Both P1 and P2 are correct w.r.t. the given pre- and postcondition in standard Hoare logic. However, only P1 is access secure w.r.t. the given pre- and postcondition, while P2 is not: the given precondition is not necessary for the postcondition to hold after executing P2, as P2 sets acc to true in any case.
@@ 2.2. Access Security for Bitcoin
In this context, the focus is on the access security of the locking script: In any successful run of the unlocking and then the locking script, the unlocking script is required to finish in a state which fulfills a suitable precondition, see Fig. 1. For a typical locking script like the standard locking script $\mathtt{scriptPubKey}_{\mathtt{pubKeyHash}}$ (bitcoin:wiki), a typical precondition would express that an address and a signature are provided, where the address hashes to a value stored in the locking script, and the signature matches the address.
direction of executionunlocking scriptlocking scriptpreconditionpostcondition
@@ 2.3. Access Security for While Loops
Suitable pre- and postconditions for certifying the access security of this program should reflect the requirement that access should only be granted if a given passkey $\mathtt{p}$ is included in a list of stored passkeys $\mathtt{L}$. A suitable postcondition is given by $\mathtt{acc}=\mathtt{true}$. A suitable precondition for when to give access is that $\mathtt{p}$ occurs in $\mathtt{L}$, which may be expressed as $\exists j\leq\mathtt{lh}(\mathtt{L})(\mathtt{el}(j,\mathtt{L})=\mathtt{p})$. A suitable invariant for the while loop that reflects the intended behaviour of the program may be $\mathtt{acc}=\mathtt{true}\ \lor\ \exists j\leq\mathtt{lh}(\mathtt{L})(j\geq i\land\mathtt{el}(j,\mathtt{L})=\mathtt{p})$ where $i$ denotes the current index of the while loop. The above loop invariant reflects the intention that within the while loop, to reach the postcondition it is necessary that either the postcondition has been reached already ($\mathtt{acc}=\mathtt{true}$), or that a matching passkey in the list $\mathtt{L}$ will be found in the remaining steps of the loop ($\exists j\leq\mathtt{lh}(\mathtt{L})(j\geq i\land\mathtt{el}(j,\mathtt{L})=\mathtt{p})$).
@@ 3. Access Hoare Triples
Hoare logic is using Hoare triples to describe how the execution of code impacts on the state of the machine. The Hoare triple $\{P\}\,C\,\{Q\}$ consists of assertions $P$ and $Q$, and a piece of code $C$. $P$ is called the precondition, $Q$ the postcondition. $\{P\}\,C\,\{Q\}$ expresses that when $P$ is true and $C$ is executed and terminates then $Q$ is true. In other words, $P$ is a sufficient condition that $Q$ holds after successfully executing $C$. It can be expressed more formally by quantifying over all possible states:
@@
We are interested in the opposite direction, that $P$ is a necessary condition for $Q$ to hold after executing $C$. We again use a triple $\langle P\rangle\,C\,\langle Q\rangle$, which we denote access Hoare triple, consisting again of a precondition $P$, a postcondition $Q$, and a program $C$. However, now $\langle P\rangle\,C\,\langle Q\rangle$ expresses that when the execution of $C$ results in $Q$ being true, then $P$ must have been true before, which can be expressed more formally as
@@
At this point we could introduce a calculus for access Hoare logic indirectly as negated Hoare logic in the sense of Theorem 3.1. However, it is of advantage to have a direct calculus for access Hoare logic. For example, a calculus for access Hoare logic would form the basis for systems that can verify access security in the sense of access Hoare logic using tools like theorem provers, similar to SPARK (SPARKAda:Homepage) and Dafny (dafny:Homepage) for verifying correctness based on standard Hoare logic. A reduction from access Hoare logic to standard Hoare logic, as outlined above, would introduce the contraposition of implications in verification proofs, i.e. $\neg B\rightarrow\neg A$ instead of $A\rightarrow B$. [^1] In more detail, a proof of an access Hoare triple $\langle P\rangle\,C\,\langle Q\rangle$ will usually involve subproofs of access Hoare triples of the form $\langle P^{\prime}\rangle\,C^{\prime}\,\langle Q^{\prime}\rangle$ and $\langle P^{\prime\prime}\rangle\,C^{\prime\prime}\,\langle Q^{\prime\prime}\rangle$, for certain subprograms $C^{\prime}$ and $C^{\prime\prime}$, where $C^{\prime}$ is immediately followed by $C^{\prime\prime}$. In such a situation, arguing access Hoare logic style from post- to preconditions requires proving the implication $P^{\prime\prime}\rightarrow Q^{\prime}$ from the precondition $P^{\prime\prime}$ of the second subproof to the postcondition $Q^{\prime}$ of the first. If we consider this situation by transformation to Hoare logic, we are faced with $\{\neg P\}\,C\,\{\neg Q\}$ and subproofs of $\{\neg P^{\prime}\}\,C^{\prime}\,\{\neg Q^{\prime}\}$ and $\{\neg P^{\prime\prime}\}\,C^{\prime\prime}\,\{\neg Q^{\prime\prime}\}$, now having to prove the implication $\neg Q^{\prime}\rightarrow\neg P^{\prime\prime}$ while arguing Hoare logic style from pre- to postconditions. For example, a straightforward implication such as $x=2\rightarrow x+2=4$ may translate into its contrapositive $\neg(x+2=4)\rightarrow\neg(x=2)$. While logically equivalent, [^2] In classical logic the latter is more difficult to understand and deal with in verification tasks. In general, this may even break verifiability. [^3] SPARK (SPARKAda:Homepage), which supports Hoare logic, tries to prove formulas generated by a verification process using an SMT solver. If that fails, SPARK offers to use the interactive theorem prover Coq/Rocq (sparkAda:Doc:Coq) instead. Another option when facing unprovability in SPARK, is to analyze formulas manually in Why3 (why3:homepage; adacore:SPARKdoc:ArchitectureQualityAssurance:Why3), in order to find a reason for unprovability. In the last two cases, the syntactical form of formulas matters: For example, Coq/Rocq is based on intuitionistic logic, hence transformation of formulas need to respect intuitionistic logic.
@@ 4. The Calculus
We define the calculus for $\langle P\rangle\,C\,\langle Q\rangle$, which we denote access Hoare logic, following that for Hoare logic (hoare:69), and show that it is sound with respect to the interpretation of access Hoare triples as given in (2) in the previous section. The development will focus on the $\mathtt{while}$ language similar to (hoare:69), a standard imperative language. We will show completeness in Section 6.
@@ 4.3. Consequence rule
The consequence rule is the main deviation from Hoare’s original calculus, reflecting that we are interested in expressing that the precondition is necessary for the postcondition. Thus, contrary to standard Hoare logic, we allow to weaken the precondition, and to strengthen the postcondition.
@@ 4.6. While rule
Similarly to the conditional rule, the precondition of the premise changes from conjunction to implication compared to standard Hoare logic. Furthermore, we weaken the postcondition to only condition reachable states, namely on those satisfying $\neg B$ – this is needed to be able to prove completeness in Section 6.
@@
5. Weakest and Strongest Preconditions
The aim of this section is to show a connection between Hoare triples and access Hoare triples, namely that modulo termination, weakest preconditions for Hoare triples, and strongest preconditions for access Hoare triples coincide. We start by repeating standard definitions of weakest preconditions for Hoare triples, and provide a definition of strongest preconditions for access Hoare triples. We then give explicit characterizations of weakest preconditions for Hoare triples and strongest preconditions for access Hoare triples. We conclude by showing that, for a fixed program and post condition, the strongest precondition for Hoare triples equals to the intersection of the weakest precondition for Hoare Triples with the set of states on which the program terminates.
Definition 5.1 (Weakest precondition for Hoare triples).
Given a program $C$ and a postcondition $Q$, a weakest precondition is a predicate $P^{\prime}$ such that

### v3

@@ Abstract.
Following Hoare’s seminal invention, now called Hoare logic, to reason about correctness of computer programs, we advocate a related but fundamentally different approach to reason about access security of computer programs such as access control. We define the formalism, which we denote access Hoare logic, and present examples which demonstrate its usefulness and fundamental difference to Hoare logic. We prove soundness and completeness of access Hoare logic, and provide a link between access Hoare logic and standard Hoare logic. We also demonstrate a fundamental difference of access Hoare logic to other approaches, in particular incorrectness logic.
@@
copyright: acmcopyrightjournalyear: XXXXdoi: XXXXXXX.XXXXXXXjournal: TOCLjournalvolume: XXjournalnumber: Xarticle: 111publicationmonth: 8ccs: Theory of computation Hoare logicccs: Theory of computation Logic and verificationccs: Software and its engineering Software verificationccs: Security and privacy Logic and verificationccs: Security and privacy Software security engineering
@@ 1. Introduction
The correctness of programs has been a concern early on in the development of computers. The seminal contribution by Tony Hoare (Hoare, 1969) allows to reason about programs instruction by instruction using pre- and post-conditions. It allows to ascertain that if the pre-condition is satisfied before execution, and program execution terminates, then the post-condition will be satisfied.
@@
This form of ascertaining the correctness of programs has been and still is very successful. It is at the heart of several tools that are used in academia and industry to specify and verify software systems, such as SPARK (Adacore, 2025b) and Dafny (Dafny Development Team, 2025).
@@
However, there are security properties that do not naturally fall within the framework of Hoare logic. Access control is the property of restricting access to a resource. It is a well-established area of computer security that includes identification, authorization, authentication, access approval, and auditing (Gollmann, 2011), where access approval is the process of deciding whether to grant or reject an access request. Access control has a broad range of applications, ranging from large distributed systems to local systems such as electronic keys. More recently, it gained importance in the context of blockchain and cryptocurrencies: Smart contracts are often written to control access to digital assets, such as cryptocurrencies, tokens, or other forms of digital assets managed by blockchain systems, thereby combining several properties of general access control. We will use the term access security to speak about such properties in general.
@@
For a program to be access secure, we need to guarantee that the final state (e.g., access to a hotel room via electronic keys) will only be granted if a given set of conditions has been met before execution (e.g., possession of valid access codes). That is, for access security we are interested in the reverse direction of Hoare logic, which we will denote access Hoare logic (aHl), where we will reason from post-conditions to pre-conditions. While normal Hoare logic expresses a pre-condition to be sufficient for the post-condition to hold after executing a program, access Hoare logic requires the pre-condition to be necessary. More precisely, access Hoare logic expresses that if the program starting in state $s$ terminates in a state $s^{\prime}$ which fulfills the post-condition, then it is necessary that state $s$ fulfills the pre-condition. To make the difference between access Hoare logic and standard Hoare logic more apparent, we observe that while False is always a valid pre-condition in normal Hoare logic, in case of access Hoare logic this is given by True.
@@
In this note, we define access Hoare triples and access Hoare logic, and use it to describe access security for examples. We show that access Hoare logic is sound and complete for access Hoare triples under a general, semantical interpretation. We provide a link between Hoare logic and access Hoare logic in that weakest pre-conditions for usual Hoare logic correspond to strongest pre-conditions for access Hoare logic on states where the program terminates (see Section 5). We also demonstrate a fundamental difference of access Hoare logic to incorrectness logic (O’Hearn, 2019) and related approaches: while incorrectness logic provides a complementary view to standard Hoare logic in the context of forward transformers, access Hoare logic plays a similar role in the context of backwards transformers (see Section 8).
@@
Alhabadi et al. (Alhabardi et al., 2022) have recently discovered that weakest pre-conditions for Hoare logic can be used to express access security for Bitcoin scripts. This paper expands on those insights and provides a reason why that is the case.
@@ 2. Examples
We demonstrate our point of view by considering three examples: The first involves electronic keys, the second owning and transferring Bitcoins, and the third granting access depending on the presence of a key in a list of keys. We will revisit the examples in Section 7 to formally analyze their access security after introducing our framework.
@@ 2.1. Access Security for Electronic Keys
Our first example is a digital key system in a hotel that uses cards to open doors, which was specified and verified by Jackson (Jackson, 2012) using Alloy, and by Nipkow (Nipkow, 2006) using Isabelle/HOL. In this system, the battery-powered door locks are not connected to a network. Instead, their state is controlled by cards that are inserted into them, or, in wireless systems, moved close to them. A lock holds an internal electronic key $\mathtt{dk}$ which controls its state. A card holds two electronic keys, $\mathtt{ck1}$ for the first key and $\mathtt{ck2}$ for the second. When a new guest receives a card (for example, from the hotel reception), it will contain the key corresponding to a previous guest as $\mathtt{ck1}$, and a new fresh key as $\mathtt{ck2}$. An access control program for the door lock suitable for use in such an electronic hotel key system should give access to the room in an access-secure way, by setting a variable $\mathtt{acc}$ to true: When the guest uses their card for the first time, access will be granted if the previous guest’s key is stored as the door key ($\mathtt{dk}{}=\mathtt{ck1}{}$), in which case the door key is updated to the new key $\mathtt{ck2}$, preventing the previous guest from gaining access to the room thereafter. The next time the new guest swipes their card, access will be granted because the second key on their card, $\mathtt{ck2}$, now matches the door key ($\mathtt{dk}{}=\mathtt{ck2}{}$).
For this example, a pre-condition for verifying access security is given by $\mathtt{ck1}=\mathtt{dk}\lor\mathtt{ck2}=\mathtt{dk}$, expressing that either of the first key or the second key is the door key. A post-condition is $\mathtt{acc}=\mathtt{true}$, expressing that access is granted.
@@
Both P1 and P2 are correct w.r.t. the given pre- and post-condition in standard Hoare logic. However, only P1 is access secure w.r.t. the given pre- and post-condition, while P2 is not: the given pre-condition is not necessary for the post-condition to hold after executing P2, as P2 sets acc to true in any case. See Section 7.1 for further details.
@@ 2.2. Access Security for Bitcoin
In this context, the focus is on the access security of the locking script: In any successful run of the unlocking and then the locking script, the unlocking script is required to finish in a state which fulfills a suitable pre-condition, see Fig. 1. For a typical locking script like the locking script $\mathtt{scriptPubKey}_{\mathtt{P2PKH}}$ of the standard pay-to-public-key-hash (P2PKH) script (Bitcoin Community, 2025), a typical pre-condition would express that an address and a signature are provided, where the address hashes to a value stored in the locking script, and the signature matches the address. Further details are given in Section 7.2.
direction of executionunlocking scriptlocking scriptpre-conditionpost-condition
@@ 2.3. Access Security for While Loops
Suitable pre- and post-conditions for certifying the access security of this program should reflect the requirement that access should only be granted if a given passkey $\mathtt{p}$ is included in a list of stored passkeys $\mathtt{L}$. A suitable post-condition is given by $\mathtt{acc}=\mathtt{true}$. A suitable pre-condition for when to give access is that $\mathtt{p}$ occurs in $\mathtt{L}$, which may be expressed as $\exists j{\leq}\mathtt{lh}(\mathtt{L})(\mathtt{el}(j,\mathtt{L})=\mathtt{p})$. A suitable invariant for the while loop that reflects the intended behaviour of the program may be $\mathtt{acc}=\mathtt{true}\ \lor\ \exists j{\leq}\mathtt{lh}(\mathtt{L})(j\geq i\land\mathtt{el}(j,\mathtt{L})=\mathtt{p})$ where $i$ denotes the current index of the while loop. The above loop invariant reflects the intention that within the while loop, to reach the post-condition it is necessary that either the post-condition has been reached already ($\mathtt{acc}=\mathtt{true}$), or that a matching passkey in the list $\mathtt{L}$ will be found in the remaining steps of the loop ($\exists j{\leq}\mathtt{lh}(\mathtt{L})(j\geq i\land\mathtt{el}(j,\mathtt{L})=\mathtt{p})$). A detailed account of the access security of CheckList is given in Section 7.3.
@@ 3. Access Hoare Triples
Following Hoare (Hoare, 1969), we take assertions to be semantical, that is, with no fixed syntax, and no restrictions on the expressions in the language. The reason for this approach is that we want to focus on the general properties of access Hoare logic and their comparison to Hoare logic and related frameworks.
Hoare logic is using Hoare triples to describe how the execution of code impacts on the state of the machine. The Hoare triple $\{P\}\,C\,\{Q\}$ consists of assertions $P$ and $Q$, and a piece of code $C$. $P$ is called the pre-condition, $Q$ the post-condition. $\{P\}\,C\,\{Q\}$ expresses that when $P$ is true and $C$ is executed and terminates then $Q$ is true. In other words, $P$ is a sufficient condition that $Q$ holds after successfully executing $C$. It can be expressed more formally by quantifying over all possible states:
@@
We are interested in the opposite direction, that $P$ is a necessary condition for $Q$ to hold after executing $C$. We again use a triple $\langle P\rangle\,C\,\langle Q\rangle$, which we denote access Hoare triple, consisting again of a pre-condition $P$, a post-condition $Q$, and a program $C$. However, now $\langle P\rangle\,C\,\langle Q\rangle$ expresses that when the execution of $C$ results in $Q$ being true, then $P$ must have been true before, which can be expressed more formally as
@@
At this point we could introduce a calculus for access Hoare logic indirectly as negated Hoare logic in the sense of Theorem 3.1. However, it is of advantage to have a direct calculus for access Hoare logic. For example, a calculus for access Hoare logic would form the basis for systems that can verify access security in the sense of access Hoare logic using tools like theorem provers, similar to SPARK (Adacore, 2025b) and Dafny (Dafny Development Team, 2025) for verifying correctness based on standard Hoare logic. A reduction from access Hoare logic to standard Hoare logic, as outlined above, would introduce the contraposition of implications in verification proofs, i.e. $\neg B\rightarrow\neg A$ instead of $A\rightarrow B$. [^1] In more detail, a proof of an access Hoare triple $\langle P\rangle\,C\,\langle Q\rangle$ will usually involve subproofs of access Hoare triples of the form $\langle P^{\prime}\rangle\,C^{\prime}\,\langle Q^{\prime}\rangle$ and $\langle P^{\prime\prime}\rangle\,C^{\prime\prime}\,\langle Q^{\prime\prime}\rangle$, for certain subprograms $C^{\prime}$ and $C^{\prime\prime}$, where $C^{\prime}$ is immediately followed by $C^{\prime\prime}$. In such a situation, arguing access Hoare logic style from post- to pre-conditions requires proving the implication $P^{\prime\prime}\rightarrow Q^{\prime}$ from the pre-condition $P^{\prime\prime}$ of the second subproof to the post-condition $Q^{\prime}$ of the first. If we consider this situation by transformation to Hoare logic, we are faced with $\{\neg P\}\,C\,\{\neg Q\}$ and subproofs of $\{\neg P^{\prime}\}\,C^{\prime}\,\{\neg Q^{\prime}\}$ and $\{\neg P^{\prime\prime}\}\,C^{\prime\prime}\,\{\neg Q^{\prime\prime}\}$, now having to prove the implication $\neg Q^{\prime}\rightarrow\neg P^{\prime\prime}$ while arguing Hoare logic style from pre- to post-conditions. For example, a straightforward implication such as $x=2\rightarrow x+2=4$ may translate into its contrapositive $\neg(x+2=4)\rightarrow\neg(x=2)$. While logically equivalent, [^2] In classical logic the latter is more difficult to understand and deal with in verification tasks. In general, this may even break verifiability. [^3] SPARK (Adacore, 2025b), which supports Hoare logic, tries to prove formulas generated by a verification process using an SMT solver. If that fails, SPARK offers to use the interactive theorem prover Coq/Rocq (Adacore, 2025a) instead. Another option when facing unprovability in SPARK, is to analyze formulas manually in Why3 (Why3, 2025; Adacore, 2025c), in order to find a reason for unprovability. In the last two cases, the syntactical form of formulas matters: For example, Coq/Rocq is based on intuitionistic logic, hence transformation of formulas need to respect intuitionistic logic.
@@ 4. The Calculus
We define the calculus for $\langle P\rangle\,C\,\langle Q\rangle$, which we denote access Hoare logic, following that for Hoare logic (Hoare, 1969), and show that it is sound with respect to the interpretation of access Hoare triples as given in (2) in the previous section. The development will focus on the $\mathtt{while}$ language similar to (Hoare, 1969), a standard imperative language. We will show completeness in Section 6.
@@ 4.3. Consequence rule
The consequence rule is the main deviation from Hoare’s original calculus, reflecting that we are interested in expressing that the pre-condition is necessary for the post-condition. Thus, contrary to standard Hoare logic, we allow to weaken the pre-condition, and to strengthen the post-condition.
@@ 4.6. While rule
Similarly to the conditional rule, the pre-condition of the premise changes from conjunction to implication compared to standard Hoare logic. Furthermore, we weaken the post-condition to only condition reachable states, namely on those satisfying $\neg B$ – this is needed to be able to prove completeness in Section 6.
@@
5. Weakest and Strongest Pre-Conditions
The aim of this section is to show a connection between Hoare triples and access Hoare triples, namely that modulo termination, weakest pre-conditions for Hoare triples, and strongest pre-conditions for access Hoare triples coincide. We start by repeating standard definitions of weakest pre-conditions for Hoare triples, and provide a definition of strongest pre-conditions for access Hoare triples. We then give explicit characterizations of weakest pre-conditions for Hoare triples and strongest pre-conditions for access Hoare triples. We conclude by showing that, for a fixed program and post condition, the strongest pre-condition for Hoare triples equals to the intersection of the weakest pre-condition for Hoare Triples with the set of states on which the program terminates.
Definition 5.1 (Weakest pre-condition for Hoare triples).
Given a program $C$ and a post-condition $Q$, a weakest pre-condition is a predicate $P^{\prime}$ such that